# Special Hamiltonian $S^{1}$-actions on symplectic 4-manifolds

### v3

@@ 5 A special Hamiltonian $S^{1}$-action on $\mathcal{W}_{n}$
Let $(M,\omega,G,h)$ be a connected and 1-connected exact special Hamiltonian $S^{1}$ space with $n+1$ fixed points, $n\in\mathbb{N}\cup\{0\}$, and $X$ the Hamiltonian vector field associated to $h$. Let $S=S_{1}\cup S_{2}\cdots\cup S_{n}$ be an $A_{n}$-string of Lagrangian spheres in $h^{-1}(0)\subset(M,\omega,G)$ associated with $(G,h)$, and $\textsf{S}:=\textsf{S}_{1}\cup\textsf{S}_{2}\cdots\cup\textsf{S}_{n}$ the $A_{n}$-string of Lagrangian spheres in $(\mathcal{W}_{n},\omega_{n},\textsf{G})$, say, induced from (5). There exists a diffeomorphism $\phi:S\to\textsf{S}$ so that $\phi(S_{i})=\textsf{S}_{i}$ for $i=1,\cdots,n$, and $\textsf{G}\circ\phi=\phi\circ G$ on S. Identify $\mathcal{W}_{n}=\cup_{i=1}^{n}T^{*}\textsf{S}_{i}=\cup_{i=1}^{n}T^{*}S_{i}$ with the plumbing of $n$ copies of the cotangent bundle $T^{*}S^{2}$ of Lagrangian $2$-spheres. By Darboux theorem we may extend $\phi$ to a symplectomorphism from a neighborhood of $S_{i}\cap S_{i+1}$ to a neighborhood of $\textsf{S}_{i}\cap\textsf{S}_{i+1}$ for each $i$. Since a diffeomorphism between two Lagrangian surfaces always extends to a symplectomorphism between their cotangent bundles, we can extend $\phi$ to a symplectomorphism $\phi:M\to\mathcal{W}_{n}$, sending $T^{*}S_{i}$ to $T^{*}\textsf{S}_{i}$ so that $\phi^{*}\omega_{n}=\omega$.
Proof of Theorem 1.6(i).
@@
Recall that the $n=0$ case follows from Proposition 3.1 (ii), so we may assume that $n\geq 1$. With $\phi$ in the proof of Theorem 1.4 we may identify $(M,\omega,G,h)$ with $(\mathcal{W}_{n},\omega_{n},G,h)$, so that $G$ acts on $(\mathcal{W}_{n},\omega_{n}=d\alpha_{n})$ as an $S^{1}$-group of Hamiltonian diffeomorphisms $G_{\theta}\in Ham(\mathcal{W}_{n},\omega_{n})$, $G_{0}=id$. Both $G$ and G actions preserve $\textsf{S}=\textsf{S}_{1}\cup\cdots\cup\textsf{S}_{n}$ with $Fix(G)=Fix(\textsf{G})=\{p_{0},p_{1},\cdots,p_{n}\}$, $p_{0}\in\textsf{S}_{1}$, $p_{i}=\textsf{S}_{i}\cap\textsf{S}_{i+1}$ for $i=1,\cdots,n-1$, and $p_{n}\in\textsf{S}_{n}$. Since both $G$ and G acts on each $\textsf{S}_{i}$ as rotations with $p_{i-1}$ and $p_{i}$ fixed, by modifying $\phi$ if necessary we may assume that $G_{\theta}=\textsf{G}_{\theta}$ on S for all $\theta$. The Euclidean metric $g_{0}$ on $\mathbb{C}^{3}$ restricts to a G-invariant, $\omega_{n}$-compatible Riemannian metric on $\mathcal{W}_{n}$. Also let $g$ denote a $\omega_{n}$-compatible metric on $\mathcal{W}_{n}$ which is invariant under the $G$-action. We may assume that $g=g_{0}$ on $T_{\textsf{S}_{i}}\mathcal{W}_{n}$ the tangent bundle of $\mathcal{W}_{n}$ over $\textsf{S}_{i}$ for all $i$. Let $N(\textsf{S}_{i})$ denote the normal bundle of $\textsf{S}_{i}\subset\mathcal{W}_{n}$ determined by the metric $g_{0}$, which can also be identified with the cotangent bundle $T^{*}\textsf{S}_{i}$. Note that
$N_{p}(\textsf{S}_{i})=T_{p}\textsf{S}_{j},\qquad T_{p}\textsf{S}_{i}=N_{p}(\textsf{S}_{j})\ \text{ at $p=\textsf{S}_{i}\cap\textsf{S}_{j}$ when $|i-j|=1$}.$ (6)
@@
Recall that the G-action on $\mathcal{W}_{n}$ is linear. Since $G_{\theta}=\textsf{G}_{\theta}$ on $\textsf{S}_{i}$, the linearized $G$-action $G^{L}:=G_{*}$ on $N(S_{i})$ acts on $N(\textsf{S}_{i})$ in the same way as $\textsf{G}=\textsf{G}_{*}$. Denote $x_{\theta}:=G_{\theta}(x)=\textsf{G}_{\theta}(x)$ and $v_{\theta}:=(G_{\theta})_{*}(v)=\textsf{G}_{\theta}(v)\in N_{x_{\theta}}(\textsf{S}_{i})$ for $x\in(\textsf{S}_{i})$, $v\in N_{x}(\textsf{S}_{i})$. We have for $v\in N_{x}(\textsf{S}_{i})$
$(G_{\theta})_{*}=\textsf{G}_{\theta}:N_{x}(\textsf{S}_{i})\to N_{x_{\theta}}(\textsf{S}_{i}),\quad v\to v_{\theta}=\textsf{G}_{\theta}(v).$
At $p_{i}\in Fix(G)$ we have
$(G_{\theta})_{*}=\textsf{G}_{\theta}:T_{p_{i}}\mathcal{W}_{n}\to T_{p_{i}}\mathcal{W}_{n}$
@@
If $n>1$ then at each $p_{i}\in Fix(G)$ with $1\leq i<n$ we also include the exponential map
@@
relative to any $G$-invariant $\omega_{n}$-compatible Riemannian metric $g$: for $v\in T_{p_{i}}\mathcal{W}_{n}$
@@
Since $^{L}:=G_{*}=\textsf{G}$ G and $g=g_{0}$ on $T_{\textsf{S}}\mathcal{W}_{n}$, $\gamma_{v}$ and $\gamma^{0}_{v}$ are tangent at the base point $x$ of $v$. For $v\in N(\textsf{S})$ denote $v_{\theta}:=G^{L}_{\theta}(v)$, $v_{0}=v$, then
$G^{L}_{\theta}(tv)=tv_{\theta},\quad G^{L}_{\theta_{1}+\theta_{2}}(v)=G^{L}_{\theta_{1}}G^{L}_{\theta_{2}}(v)=G^{L}_{\theta_{1}}(v_{\theta_{2}})=v_{\theta_{1}+\theta_{2}}.$
Since $g$ is $G$-invariant, $G_{\theta}$ maps $g$-geodesics to $g$-geodesics, preserving the arc length. Also, $G_{\theta}=\textsf{G}_{\theta}$ when restricted to S, we have $G_{\theta}(\gamma_{v})=\gamma_{v_{\theta}}$, and for $v\in N(\textsf{G})$ based at $x\in\textsf{S}$, the map $\EuFrak{e}_{\textsf{S}}$ takes $\gamma^{0}_{v}(1)$ to $\gamma_{v}(1)$, and $\gamma^{0}_{v_{\theta}}(1)$ to $\gamma_{v_{\theta}}(1)$.
The domain $U:=\EuFrak{e}_{\textsf{S}}(V)$ is $G$-invariant since $V$ is $^{L}=\textsf{G}$ G-invariant and $g=g^{0}$ on $V\subset N(\textsf{S})$. Also we may identify $(V\subset N(\textsf{S}),\omega_{n})$ with its image $(\EuFrak{e}_{\textsf{S}}(V)=U,\omega_{n})$. Then on $U$
$G=\EuFrak{e}_{\textsf{S}}\circ\textsf{G}\circ(\EuFrak{e}_{\textsf{S}})^{-1}.$
@@
$\omega^{\prime}:=((\EuFrak{e}_{\textsf{S}})^{-1})^{*}\omega_{n}=((\EuFrak{e}_{\textsf{S}})^{-1})^{*}\sum_{j=1}^{3}dx_{j}\wedge dy_{j}=\sum_{j=1}^{3}dx^{\prime}_{j}\wedge dy^{\prime}_{j},$
[…]
@@
is the the corresponding Hamiltonian vector field and Hamiltonian function associated to the $G$-action on $U$. We may extend $\EuFrak{e}_{\textsf{S}}$ to a diffeomorphism $\EuFrak{e}^{\prime}_{\textsf{S}}:\mathcal{W}_{n}\to\mathcal{W}_{n}$ with $G$-invariant compact support containing $U$. Let $G^{\prime}:=\EuFrak{e}^{\prime}_{\textsf{S}}\circ\textsf{G}\circ(\EuFrak{e}^{\prime}_{\textsf{S}})^{-1}=p$. $G^{\prime}$ acts as a Hamiltonian $S^{1}$-action $\mathcal{W}_{n}$, with $G^{\prime}=G$ on $U$ and outside the compact support of $\EuFrak{e}^{\prime}_{\textsf{S}}$.
Note that $(\EuFrak{e}_{\textsf{S}})_{*}:=d\EuFrak{e}_{\textsf{S}}=id$ when restricted to S, so $\omega^{\prime}=\omega_{n}$ on S. Both $\omega^{\prime}=d\alpha^{\prime}$ and $\omega_{n}=d\alpha_{n}$ are $G$-invariant exact symplectic forms. Up to averaging via the $G$-action we may assume that both primitive 1-forms $\alpha_{n}$ and $\alpha^{\prime}$ are $G$-invariant. Then similar to the proof of Proposition 3.1 (ii), by shrinking $U$ if necessarily, there exists a smooth isotopy $\phi_{t}:U\to M$, $t\in[0,1]$, with $\phi_{0}=id$, $\phi_{t}=id$ on S for all $t$, and $\phi_{1}^{*}\omega^{\prime}=\omega_{n}$. Therefore, any effective special Hamiltonian $S^{1}$-action on $(\mathcal{W}_{n},\omega_{n})$ is linear near S up to an equivariant smooth isotopy $\phi_{t}$ with $\phi^{-1}_{1}\circ G\circ\phi_{1}=\textsf{G}$ on $U$. ∎
In the following we will show that on $(\mathcal{W}_{n},\omega_{n})$ with $n=0,1$, any semi-free Hamiltonian $S^{1}$-action, which we denote it as $G$, up to conjugations is symplectically equivariant to the G-action on any prescribed relatively compact domain containing (i) the fixed point of the G-action if $n=0$, and (ii) a G-invariant Lagrangian sphere if $n=1$.
Proof of Theorem 1.6(ii)(iii).
Case 1: $n=0$. We denote $q=(x,y)$ for $q\in(\mathbb{C}^{2},\omega=\sum_{j=1}^{2}dx_{j}\wedge dy_{j})$, where $x=(x_{1},x_{2})$, $y=(y_{1},y_{2})$, also $|x|=\sqrt{x_{1}^{2}+x_{2}^{2}}$, $|y|=\sqrt{y_{1}^{2}+y_{2}^{2}}$.
@@
Consider The Liouville vector field $Y:=\frac{1}{2}\sum_{j=1}^{2}x_{j}\partial_{x_{j}}+y_{j}\partial_{y_{j}}$ on $\mathbb{C}^{2}$. The differential $\iota_{*}:T\mathbb{C}^{2}\to T\mathcal{W}_{0}$ induces a Liouville vector field $Y_{0}$ on $\mathcal{W}_{0}$ given by $Y_{0}:=\iota_{*}Y$. Both $Y$ and $Y_{0}$ are G-invariant. Recall from Proposition 3.1 $G$ and G are symplectically equivariant when restricted to some open neighborhood $U$ of the fixed point. Since $\mathbb{C}_{2}$ and $\mathcal{W}_{0}$ are symplectomorphic G-spaces, and any semi-free special Hamiltonian $S^{1}$-action on $\mathcal{W}_{0}$ descends to one in $\mathbb{C}^{2}$, so it suffices to compare $G$ and G in $(\mathbb{C}^{2},\omega)$. Then by Proposition 3.1 (ii) up to conjugation $G$-action is symplectically equivariant to G-action on some open neighborhood $U_{0}$ of $0\in\mathbb{C}^{2}$, i.e., there exists a diffeomorphism $\EuFrak{e}_{0}:U_{0}\to U$ from a G-invariant open neighborhood of $0\in\mathbb{C}^{2}$ onto a $G$-invariant open neighborhood $U$ of $0\in\mathbb{C}^{2}$, $\EuFrak{e}_{0}(0)=0\in\mathbb{C}^{2}$, and
$G_{\theta}=\EuFrak{e}_{0}\circ\textsf{G}_{\theta}\circ\EuFrak{e}_{0}^{-1}\quad\text{on $U$},\quad\forall\theta,$
@@
$\textsf{G}_{\theta}=\EuFrak{e}_{0}^{-1}\circ G_{\theta}\circ\EuFrak{e}_{0}\quad\text{on $U_{0}$},\quad\forall\theta.$
@@
$\begin{split}(\Psi_{r}^{-1}\circ G_{\theta}\circ\Psi_{r})(x,y)&=\psi_{t_{r}}^{-1}\circ\EuFrak{e}_{0}^{-1}\circ G_{\theta}\circ\EuFrak{e}_{0}\circ\psi_{t_{r}}(x,y)\\ &=\psi_{t_{r}}^{-1}\circ\EuFrak{e}_{0}^{-1}\circ G_{\theta}\circ\EuFrak{e}_{0}(e^{t_{r}/2}x,e^{t_{r}/2}y)\\ &=\psi_{t_{r}}^{-1}\circ\textsf{G}_{\theta}(e^{t_{r}/2}x,e^{t_{r}/2}y)\\ &=\psi_{t_{r}}^{-1}(e^{t_{r}/2}x_{\theta},e^{t_{r}/2}y_{\theta})\\ &=(x_{\theta},y_{\theta})=\textsf{G}_{\theta}(x,y)\end{split}$ (7)
@@
$\begin{split}x_{\theta}&:=(x_{1}\cos\theta+x_{2}\sin\theta,-x_{1}\sin\theta+x_{2}\cos\theta),\\ y_{\theta}&:=(y_{1}\cos\theta+y_{2}\sin\theta,-y_{1}\sin\theta+y_{2}\cos\theta).\end{split}$ (8)
@@
$\Psi_{r}^{-1}\circ G_{\theta}\circ\Psi_{r}=\textsf{G}\quad\text{on }\{x^{2}+y^{2}\leq r<r_{1}\}\subset\mathbb{C}^{2}.$
By allowing $r_{1}$ and $r_{0}$ to be arbitrarily large we conclude that any semi-free special Hamiltonian $S^{1}$-action on $(\mathbb{C}^{2},\omega)$ is eventually symplectically equivariant to the standard linear action of G. This completes the proof of Case 1.
Case 2: $n=1$.  Consider $(\mathcal{W}_{1},\omega_{1})$ where $\mathcal{W}_{1}:\{z_{1}^{2}+z_{2}^{2}+z_{3}^{2}=1\}\subset\mathbb{C}^{3}$ is given by the equations
@@
$\mathcal{W}_{1}$ is indeed the cotangent bundle of the sphere $S:=\{\sum_{j=1}^{3}x_{j}^{2}=1\}\subset\mathbb{R}^{3}$ which is Lagrangian viewed as a submanifold of $(\mathcal{W}_{1},\omega_{1})$, G acts on $S$ with two fixed points $(0,0,\pm 1)$. Consider the vector field on $\mathcal{W}_{1}$:
@@
satisfies $d\alpha=\omega_{1}$, is a G-invariant primitive 1-form of $\omega_{1}$, so $Y$ is a Liouville vector field on $\mathcal{W}_{1}$ which vanishes precisely on the Lagrangian sphere $S$. Moreover, $Y$ is invariant under the G-action, which can be checked by direct computation.

### v4

@@ 5 A special Hamiltonian $S^{1}$-action on $\mathcal{W}_{n}$
Let $(M,\omega,\Phi)$ be a connected and 1-connected exact SHam1-space with $n+1$ fixed points, $n\in\mathbb{N}\cup\{0\}$, and $X$ the Hamiltonian vector field associated to $h$. Let $S=S_{1}\cup S_{2}\cdots\cup S_{n}$ be an $A_{n}$-string of Lagrangian spheres in $h^{-1}(0)\subset(M,\omega,\Phi)$ associated with $(\Phi,h)$, and $\textsf{S}:=\textsf{S}_{1}\cup\textsf{S}_{2}\cdots\cup\textsf{S}_{n}$ the $A_{n}$-string of Lagrangian spheres in $(\mathcal{W}_{n},\omega_{n},\Theta)$, say, induced from (4). There exists a diffeomorphism $\phi:S\to\textsf{S}$ so that $\phi(S_{i})=\textsf{S}_{i}$ for $i=1,\cdots,n$, and $\Theta\circ\phi=\phi\circ\Phi$ on $S$. Identify $\mathcal{W}_{n}=\cup_{i=1}^{n}T^{*}\textsf{S}_{i}=\cup_{i=1}^{n}T^{*}S_{i}$ with the plumbing of $n$ copies of the cotangent bundle $T^{*}S^{2}$ of Lagrangian $2$-spheres. By Darboux theorem we may extend $\phi$ to a symplectomorphism, also denoted as $\phi$, $\phi:U\to\phi(U)\subset\mathcal{W}_{n}$, from a neighborhood $U$ of $S_{i}\cup S_{i+1}$ to a neighborhood $\phi(U)$ of $\textsf{S}_{i}\cup\textsf{S}_{i+1}$ for each $i$. Since a diffeomorphism between two Lagrangian surfaces always extends to a symplectomorphism between their cotangent bundles, we can extend $\phi$ to a symplectomorphism $\phi:M\to\mathcal{W}_{n}$, sending $T^{*}S_{i}$ to $T^{*}\textsf{S}_{i}$ so that $\phi^{*}\omega_{n}=\omega$. $\Box$
Proof of Theorem 1.10 (i)
Recall that the $n=0$ case follows from Proposition 3.1, so we may assume that $n\geq 1$. With $\phi$ in the proof of Theorem 1.8 we may identify $(M,\omega,\Phi,h)$ with $(\mathcal{W}_{n},\omega_{n},\Phi,h)$, so that $\Phi$ acts on $(\mathcal{W}_{n},\omega_{n}=d\alpha_{n})$ as an $S^{1}$-group of Hamiltonian diffeomorphisms $\Phi_{\theta}\in Ham(\mathcal{W}_{n},\omega_{n})$, $\Phi_{0}=id$.
@@
Both $\Phi$ and $\Theta$ actions preserve $\textsf{S}=\textsf{S}_{1}\cup\cdots\cup\textsf{S}_{n}$ with $Fix(\Phi)=Fix(\Theta)=\{p_{0},p_{1},\cdots,p_{n}\}$, $p_{0}\in\textsf{S}_{1}$, $p_{i}=\textsf{S}_{i}\cap\textsf{S}_{i+1}$ for $i=1,\cdots,n-1$, and $p_{n}\in\textsf{S}_{n}$. Since both $\Phi$ and $\Theta$ act on each $\textsf{S}_{i}$ as rotations with $p_{i-1}$ and $p_{i}$ fixed, by modifying $\Phi$ if necessary we may assume that $\Phi_{\theta}=\Theta_{\theta}$ on S for all $\theta$. The Euclidean metric $g_{0}$ on $\mathbb{C}^{3}$ restricts to a $\Theta$-invariant, $\omega_{n}$-compatible Riemannian metric on $\mathcal{W}_{n}$. Also let $g$ denote a $\omega_{n}$-compatible metric on $\mathcal{W}_{n}$ which is also $\Phi$-invariant. We may assume that $g=g_{0}$ on $T_{\textsf{S}_{i}}\mathcal{W}_{n}$ the tangent bundle of $\mathcal{W}_{n}$ over $\textsf{S}_{i}$ for all $i$. Let $N(\textsf{S}_{i})$ denote the normal bundle of $\textsf{S}_{i}\subset\mathcal{W}_{n}$ determined by the metric $g_{0}$, which can also be identified with the cotangent bundle $T^{*}\textsf{S}_{i}$. Note that
(5) $N_{p}(\textsf{S}_{i})=T_{p}\textsf{S}_{j},\qquad T_{p}\textsf{S}_{i}=N_{p}(\textsf{S}_{j})\ \text{ at $p=\textsf{S}_{i}\cap\textsf{S}_{j}$ when $|i-j|=1$}.$
@@
Recall that the action of $\Theta$ on $\mathcal{W}_{n}$ is linear. Since $\Phi_{\theta}=\Theta_{\theta}$ on $\textsf{S}_{i}$, the linearized action $\Phi_{*}$ of $\Phi$ acts on $N(S_{i})$ in the same way as $\Theta=\Theta_{*}$ does. Denote $x_{\theta}:=\Phi_{\theta}(x)=\Theta_{\theta}(x)$ and $v_{\theta}:=(\Phi_{\theta})_{*}(v)=\Theta_{\theta}(v)\in N_{x_{\theta}}(\textsf{S}_{i})$ for $x\in\textsf{S}_{i}$, $v\in N_{x}(\textsf{S}_{i})$. We have for $v\in N_{x}(\textsf{S}_{i})$
$(\Phi_{\theta})_{*}=\Theta_{\theta}:N_{x}(\textsf{S}_{i})\to N_{x_{\theta}}(\textsf{S}_{i}),\quad v\to v_{\theta}=\Theta_{\theta}(v).$
At $p_{i}\in Fix(\Phi)$ we have
$(\Phi_{\theta})_{*}=\Theta_{\theta}:T_{p_{i}}\mathcal{W}_{n}\to T_{p_{i}}\mathcal{W}_{n}$
@@
If $n>1$ then at each $p_{i}\in Fix(\Phi)$ with $1\leq i<n$ we also include the exponential map
@@
relative to any $\Phi$-invariant $\omega_{n}$-compatible Riemannian metric $g$: for $v\in T_{p_{i}}\mathcal{W}_{n}$
@@
Since $\Phi_{*}=\Theta$ and $g=g_{0}$ on $T_{\textsf{S}}\mathcal{W}_{n}$, $\gamma_{v}$ and $\gamma^{0}_{v}$ are tangent at the base point $x$ of $v$. For $v\in N(\textsf{S})$ denote $v_{\theta}:=(\Phi_{\theta})_{*}(v)$, $v_{0}=v$, then
$(\Phi_{\theta})_{*}(tv)=tv_{\theta},\quad(\Phi_{\theta_{1}+\theta_{2}})_{*}(v)=(\Phi_{\theta_{1}})_{*}(\Phi_{\theta_{2}})_{*}(v)=(\Phi_{\theta_{1}})_{*}(v_{\theta_{2}})=v_{\theta_{1}+\theta_{2}}.$
As $g$ is $\Phi$-invariant, $\Phi_{\theta}$ maps $g$-geodesics to $g$-geodesics, preserving the arc length. Also, $\Phi_{\theta}=\Theta_{\theta}$ when restricted to S, we have $\Phi_{\theta}(\gamma_{v})=\gamma_{v_{\theta}}$, and for $v\in N(\textsf{S})$ based at $x\in\textsf{S}$, the map $\EuFrak{e}_{\textsf{S}}$ takes $\gamma^{0}_{v}(1)$ to $\gamma_{v}(1)$, and $\gamma^{0}_{v_{\theta}}(1)$ to $\gamma_{v_{\theta}}(1)$.
The domain $U=\EuFrak{e}_{\textsf{S}}(V)$ is $\Phi$-invariant since $V$ is $\Phi_{*}=\Theta$-invariant and $g=g^{0}$ on $V\subset N(\textsf{S})$. Also we may identify $(V\subset N(\textsf{S}),\omega_{n})$ with its image $(\EuFrak{e}_{\textsf{S}}(V)=U,\omega_{n})$. Then on $U$
$\Phi=\EuFrak{e}_{\textsf{S}}\circ\Theta\circ\EuFrak{e}_{\textsf{S}}^{-1}.$
@@
$\omega^{\prime}:=(\EuFrak{e}_{\textsf{S}}^{-1})^{*}\omega_{n}=(\EuFrak{e}_{\textsf{S}}^{-1})^{*}\sum_{j=1}^{3}dx_{j}\wedge dy_{j}=\sum_{j=1}^{3}dx^{\prime}_{j}\wedge dy^{\prime}_{j},$
where $x^{\prime}_{j}:=x_{j}\circ\EuFrak{e}_{\textsf{S}}^{-1}$ and $y^{\prime}_{j}:=y_{j}\circ\EuFrak{e}_{\textsf{S}}^{-1}$. $\omega^{\prime}$ is symplectic on $U$, invariant under $\Phi$-action, and the action of $\Phi$ on $U$ is linear with respect to $\omega^{\prime}=\sum_{j=1}^{3}dx^{\prime}_{j}\wedge dy^{\prime}_{j}$. Let the pair $(\textsf{X},\textsf{h})$ denote the Hamiltonian vector field and corresponding Hamiltonian function associated to the $\Theta$-action, then
@@
is the the corresponding Hamiltonian vector field and Hamiltonian function associated to the $\Phi$-action on $U$. We may extend $\EuFrak{e}_{\textsf{S}}$ to a diffeomorphism $\EuFrak{e}^{\prime}_{\textsf{S}}:\mathcal{W}_{n}\to\mathcal{W}_{n}$ with $\Phi$-invariant compact support containing $U$. Let
$\Phi^{\prime}:=\EuFrak{e}^{\prime}_{\textsf{S}}\circ\Theta\circ{\EuFrak{e}^{\prime}_{\textsf{S}}}^{-1}=p.$
$\Phi^{\prime}$ acts as a SHam1–action $\mathcal{W}_{n}$, with $\Phi^{\prime}=\Phi$ on $U$ and outside the compact support of $\EuFrak{e}^{\prime}_{\textsf{S}}$.
Note that $(\EuFrak{e}_{\textsf{S}})_{*}:=d\EuFrak{e}_{\textsf{S}}=id$ when restricted to S, so $\omega^{\prime}=\omega_{n}$ on S. Both $\omega^{\prime}=d\alpha^{\prime}$ and $\omega_{n}=d\alpha_{n}$ are $\Phi$-invariant exact symplectic forms. Up to averaging via the $\Phi$-action we may assume that both primitive 1-forms $\alpha_{n}$ and $\alpha^{\prime}$ are $\Phi$-invariant. Consider the smooth family of differential forms $\omega_{t}:=(1-t)\omega_{n}+t\omega^{\prime}$, $0\leq t\leq 1$. Recall that $\omega^{\prime}=\omega_{n}$ on S. Then similar to the proof of Proposition 3.1 (ii), by shrinking $U$ if necessarily, we may assume that $\omega_{t}$ are symplectic on $U$ for all $0\leq t\leq 1$, and there exists a smooth isotopy $\phi_{t}:U\to M$, $t\in[0,1]$, with $\phi_{0}=id$, $\phi_{t}=id$ on S for all $t$, and
$\phi_{1}^{*}\omega^{\prime}=\omega_{n}.$
Therefore, any effective SHam1-action on $(\mathcal{W}_{n},\omega_{n})$ is linear near S up to conjugation with a smooth isotopy $\phi_{t}$ with $\phi^{-1}_{1}\circ\Phi\circ\phi_{1}=\Theta$ on $U$. In addition $(U,\omega^{\prime},\Phi)$ and $(V,\omega_{n},\Theta)$ are symplectically equivariant via the symplectomorphism $\EuFrak{e}_{\textsf{S}}:(U,\omega^{\prime})\to(V,\omega_{n})$. $\Box$
In the following we will show that on $(\mathcal{W}_{n},\omega_{n})$ with $n=0,1$, any semi-free SHam1-action, which we denote it as $\Phi$, up to conjugations is linear and, following1.10(i), is symplectically equivariant to the $\Theta$-action on any prescribed relatively compact domain containing (i) the fixed point of the $\Theta$-action if $n=0$, and (ii) a $\Theta$-invariant Lagrangian sphere if $n=1$.
Proof of Theorem 1.10 (ii)(iii)
(ii): $n=0$. We denote $q=(x,y)$ for $q\in(\mathbb{C}^{2},\omega=\sum_{j=1}^{2}dx_{j}\wedge dy_{j})$, where $x=(x_{1},x_{2})$, $y=(y_{1},y_{2})$, also $|x|=\sqrt{x_{1}^{2}+x_{2}^{2}}$, $|y|=\sqrt{y_{1}^{2}+y_{2}^{2}}$.
@@
Consider The Liouville vector field $Y:=\frac{1}{2}\sum_{j=1}^{2}x_{j}\partial_{x_{j}}+y_{j}\partial_{y_{j}}$ on $\mathbb{C}^{2}$. The differential $\iota_{*}:T\mathbb{C}^{2}\to T\mathcal{W}_{0}$ induces a Liouville vector field $Y_{0}$ on $\mathcal{W}_{0}$ given by $Y_{0}:=\iota_{*}Y$. Both $Y$ and $Y_{0}$ are $\Theta$-invariant. Recall from Proposition 3.1 $\Phi$ and $\Theta$ are symplectically equivariant when restricted to some open neighborhood $U$ of the fixed point. Since $\mathbb{C}_{2}$ and $\mathcal{W}_{0}$ are symplectomorphic $\Theta$-spaces, and any semi-free SHam1-action on $\mathcal{W}_{0}$ descends to one in $\mathbb{C}^{2}$, so it suffices to compare $\Phi$ and $\Theta$ in $(\mathbb{C}^{2},\omega)$. Then by Proposition 3.1 (ii) up to conjugation $\Phi$-action is equivariant to $\Theta$-action and hence is linear on some open neighborhood $U_{0}$ of $0\in\mathbb{C}^{2}$, i.e., there exists a diffeomorphism $\EuFrak{e}_{0}:U_{0}\to U$ from a $\Theta$-invariant open neighborhood of $0\in\mathbb{C}^{2}$ onto a $\Phi$-invariant open neighborhood $U$ of $0\in\mathbb{C}^{2}$, $\EuFrak{e}_{0}(0)=0\in\mathbb{C}^{2}$, and
$\Phi_{\theta}=\EuFrak{e}_{0}\circ\Theta_{\theta}\circ\EuFrak{e}_{0}^{-1}\quad\text{on $U$},\quad\forall\theta,$
@@
$\Theta_{\theta}=\EuFrak{e}_{0}^{-1}\circ\Phi_{\theta}\circ\EuFrak{e}_{0}\quad\text{on $U_{0}$},\quad\forall\theta.$
@@
(6) $\begin{split}(\Psi_{r}^{-1}\circ\Phi_{\theta}\circ\Psi_{r})(x,y)&=\psi_{t_{r}}^{-1}\circ\EuFrak{e}_{0}^{-1}\circ\Phi_{\theta}\circ\EuFrak{e}_{0}\circ\psi_{t_{r}}(x,y)\\ &=\psi_{t_{r}}^{-1}\circ\EuFrak{e}_{0}^{-1}\circ\Phi_{\theta}\circ\EuFrak{e}_{0}(e^{t_{r}/2}x,e^{t_{r}/2}y)\\ &=\psi_{t_{r}}^{-1}\circ\Theta_{\theta}(e^{t_{r}/2}x,e^{t_{r}/2}y)\\ &=\psi_{t_{r}}^{-1}(e^{t_{r}/2}x_{\theta},e^{t_{r}/2}y_{\theta})\\ &=(x_{\theta},y_{\theta})=\Theta_{\theta}(x,y)\end{split}$
@@
(7) $\begin{split}x_{\theta}&:=(x_{1}\cos\theta+x_{2}\sin\theta,-x_{1}\sin\theta+x_{2}\cos\theta),\\ y_{\theta}&:=(y_{1}\cos\theta+y_{2}\sin\theta,-y_{1}\sin\theta+y_{2}\cos\theta).\end{split}$
@@
$\Psi_{r}^{-1}\circ\Phi\circ\Psi_{r}=\Theta\quad\text{on }\{(q=(x,y)\in\mathbb{C}^{2}||q|^{2}\leq r<r_{1}\}\subset\mathbb{C}^{2}.$
By allowing $r_{1}$ and $r_{0}$ to be arbitrarily large we conclude the that any semi-free SHam1-action on $(\mathbb{C}^{2},\omega)$ with $Fix_{\Phi}=\{0\}$ is linear up to conjugation and symplectically equivariant to the standard linear action of $\Theta$ on open 4-ball $B_{r}(0)=\{q\in\mathbb{C}^{2}\mid|q|^{2}<r\}$ centered at 0 with any prescribed radius $\sqrt{r}$. This completes the proof of Case 1.
(iii): $n=1$. Consider $(\mathcal{W}_{1},\omega_{1})$ where $\mathcal{W}_{1}:\{z_{1}^{2}+z_{2}^{2}+z_{3}^{2}=1\}\subset\mathbb{C}^{3}$ is given by the equations
@@
$\mathcal{W}_{1}$ is indeed the cotangent bundle of the sphere $S:=\{\sum_{j=1}^{3}x_{j}^{2}=1\}\subset\mathbb{R}^{3}$ which is Lagrangian viewed as a submanifold of $(\mathcal{W}_{1},\omega_{1})$, $\Theta$ acts on $S$ with two fixed points $(0,0,\pm 1)$. Consider the vector field on $\mathcal{W}_{1}$:
@@
satisfies $d\alpha=\omega_{1}$, is a $\Theta$-invariant primitive 1-form of $\omega_{1}$, so $Y$ is a Liouville vector field on $\mathcal{W}_{1}$ which vanishes precisely on the Lagrangian sphere $S$. Moreover, $Y$ is invariant under the $\Theta$-action, which can be checked by direct computation.